# Transfer Learning Between U.S. Presidential Elections: How Should We Learn From A 2020 Ad Campaign To Inform 2024 Ad Campaigns?

### v1

@@ C.1 Estimating the Density Ratio
Therefore, one can obtain $\widehat{w}({\bf v})$ with $n_{s}\widehat{\mathbb{P}}(S_{i}=0\mid{\bf V}_{i}={\bf v})/\{n_{t}\widehat{\mathbb{P}}(S_{i}=1\mid]{\bf V}_{i}={\bf v})\}$, where ^$\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ can be obtained from a binary regression model or classification model. For example, when ${\cal V}$ is discrete, one may estimate this probability for any ${\bf v}\in{\cal V}$ by calculating the proportion of source samples among all samples with the same covariate:
^$\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})=\dfrac{\sum_{i=1}^{n}\mathbb{I}(S_{i}=1,{\bf V}_{i}={\bf v})}{\sum_{i=1}^{n}\mathbb{I}({\bf V}_{i}={\bf v})},\ \ \widehat{w}({\bf v})=\dfrac{n_{s}}{n_{t}}\dfrac{\widehat{\mathbb{P}}(S_{i}=0\mid{\bf V}_{i}={\bf v})}{\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})}.$ (C.1)
To account for the possible imbalance between the source and target samples (i.e., $n_{s}$ and $n_{t}$ may differ a lot), we employ the offset logistic regression proposed by Zhang et al., (2023) and extended by Kallus and Mao, (2024). A direction calculation from the Bayes rule suggests:
${\textnormal{log}}\left(\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})/\mathbb{P}(S_{i}=0\mid{\bf V}_{i}={\bf v})\right)$ $=-{\textnormal{log}}(w({\bf v}))+{\textnormal{log}}\left(\mathbb{P}(S_{i}=1)/\mathbb{P}(S_{i}=0)\right),$
which motivates positing a logistic regression model on $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ with regression parameter ${\boldsymbol{\beta}}$ and offset term ${\textnormal{log}}\left(n_{s}/n_{t}\right)$:
${\textnormal{log}}\left(\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})/\mathbb{P}(S_{i}=0\mid{\bf V}_{i}={\bf v})\right)$ $={\boldsymbol{\beta}}^{\rm T}{\bf V}_{i}+{\textnormal{log}}\left(n_{s}/n_{t}\right),\quad\widehat{w}({\bf v})=\exp(-\widehat{\boldsymbol{\beta}}^{\rm T}{\bf v}).$ (C.2)
The convergence rate for the regression coefficient has been established in Zhang et al., (2023) which can be translated to the convergence rate of $w({\bf v})$ (Kallus and Mao,, 2024). The estimation can be easily implemented using the glm package in R with the argument offset.
C.2 Proof of Theorem 4.2 and Theorem A.2
@@
We prove Theorem A.2, the EIF for a general outcome under sensitivity model (A.1). It includes Theorem 4.2 as a special case for a binary outcome. To simplify notation, we suppress the dependence of the TATE on $\gamma_{a}$ and enote the expected potential outcome on the target population at treatment level $a$ as $\theta_{a}$ for $a=0,1$. We also drop the subscript $i$ and denote by ${\bf O}$ a generic random variable, which consists of $({\bf X},Y,S=1)$ for the source and $({\bf V},S=0)$ for the target. Denote by $p_{{\bf V}\mid S=1}$, $p_{{\bf X}\mid{\bf V},S=1}$, $p_{Y\mid{\bf X},A,S=1}$ the density functions of the conditional distributions of ${\bf V}\mid S$, ${\bf X}\mid{\bf V},S=1$ and $Y\mid{\bf X},A,S=1$, respectively.
@@
We start with the case where $\pi({\bf X})$ is unknown and therefore considered as a nuisance parameter. For clarify we denote its true value as $\pi_{0}({\bf X})$. For a generic observation ${\bf O}$, the log-likelihood can be written as
@@
$\mathbb{E}\{\exp(\gamma_{1}Y)Y\mid{\bf X},A=1,S=1\}=$ $\exp(\gamma_{1})\mu_{1}({\bf X}),$
$\mathbb{E}\{\exp(\gamma_{1}Y)\mid{\bf X},A=1,S=1\}=$ $\exp(\gamma_{1})\mu_{1}({\bf X})+1-\mu_{1}({\bf X}),$
$\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\}=$ $\exp(\gamma_{1})\rho_{1}({\bf V}),$
$\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}=$ $\exp(\gamma_{1})\rho_{1}({\bf V})+1-\rho_{1}({\bf V}).$
@@
$=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\exp(\gamma_{a})}{[\exp(\gamma_{a})\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}+\mu_{a}({\bf X}_{i})-\rho_{a}({\bf V}_{i})\right]$
$+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\exp(\gamma_{a})\rho_{a}({\bf V}_{i})}{\exp(\gamma_{a})\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})},$
@@
$=$ $\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\exp(\gamma_{a})}{[\exp(\gamma_{a})\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})-\widehat{\rho}_{a}({\bf V}_{i})\right]$
$+\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\dfrac{\exp(\gamma_{a})\widehat{\rho}_{a}({\bf V}_{i})}{\exp(\gamma_{a})\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})},$
@@
In particular, if $\gamma_{1}=\gamma_{0}=0$, there exists a constant $C$ such that
@@
$=$ $\mathbb{E}\{\widehat{\varphi}_{1}({\bf O}_{i})\}-\mathbb{E}\left[\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})}{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right]$
$=$ $\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\exp(\gamma_{1})}{[\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\dfrac{A}{\widehat{\pi}({\bf X}_{i})}\{\mu_{1}({\bf X}_{i})-\widehat{\mu}_{1}({\bf X}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\exp(\gamma_{1})}{[\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})]^{2}}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\exp(\gamma_{1})}{[\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\dfrac{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})}{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})}{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right]$
$=$ $\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)\widehat{\pi}({\bf X}_{i})}\dfrac{\exp(\gamma_{1})}{[\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\left\{\dfrac{\mathbb{P}(S_{i}=1)-\widehat{\mathbb{P}}(S_{i}=1)}{\widehat{\mathbb{P}}(S_{i}=1)\mathbb{P}(S_{i}=1)}+\dfrac{1}{\mathbb{P}(S_{i}=1)}\right\}S_{i}\widehat{w}({\bf V}_{i})\dfrac{\exp(\gamma_{1})}{[\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left(\dfrac{(1-S_{i})\exp(\gamma_{1})\left[\widehat{\rho}_{1}({\bf V}_{i})\{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}-\rho_{1}({\bf V}_{i})\{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\right]}{\mathbb{P}(S_{i}=0)\widehat{\mathbb{P}}(S_{i}=0)\{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right)$
@@
$-\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\exp(\gamma_{1})}{[\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{(1-S_{i})}{\mathbb{P}(S_{i}=0)}\dfrac{\exp(\gamma_{1})}{\{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}\right]$
@@
$+\mathbb{E}\left[\dfrac{S_{i}}{\mathbb{P}(S_{i}=1)}\dfrac{\exp(\gamma_{1})\{1-\exp(\gamma_{1})\}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}\left\{\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\right\}}{\{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}^{2}\{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right]$
$+\mathbb{E}\left[\dfrac{Sw({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\exp(\gamma_{1})\{1-\exp(\gamma_{1})\}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}^{2}}{\{\exp(\gamma_{1})\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}^{2}\{\exp(\gamma_{1})\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right]$
@@
When $\gamma_{1}=0$, following the same procedure and using the fact that $\exp(\gamma_{1})=1$, we have

### v2

@@ C.2 Estimating the Density Ratio
and estimate $w({\bf v})$ by estimating $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ with a binary classifier and estimating $\mathbb{P}(S_{i}=1)$ as $n_{s}/n$; see Kallus and Mao, (2024) and Zeng et al., (2023). For example, when ${\cal V}$ is discrete, one may estimate this probability for any ${\bf v}\in{\cal V}$ by calculating the proportion of source samples among all samples with the same covariate:
^$\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})=\dfrac{\sum_{i=1}^{n}\mathds{1}(S_{i}=1,{\bf V}_{i}={\bf v})}{\sum_{i=1}^{n}\mathds{1}({\bf V}_{i}={\bf v})},\ \ \widehat{w}({\bf v})=\dfrac{n_{s}}{n_{t}}\dfrac{\widehat{\mathbb{P}}(S_{i}=0\mid{\bf V}_{i}={\bf v})}{\widehat{\mathbb{P}}(S_{i}=1\mid{\bf V}_{i}={\bf v})}.$ (C.2)
Equation (C.1) also reveals the necessity of having a sufficiently large target sample (i.e., the second part of Assumption 2.2). Intuitively, a substantially small target sample will make estimation of $\mathbb{P}(S_{i}=1\mid{\bf V}_{i}={\bf v})$ challenging due to class imbalance. Also, when $\mathbb{P}(S_{i}=0)$ is close to zero, $w({\bf v})$ can be large in magnitude, which will generally increase the bias and variance of the estimated TATE.
The second category is to use principles behind covariate balance to estimate $w({\bf v})$. Specifically, $w$ serves as a balancing score between the source and the target population, i.e.,
$\mathbb{E}\{f({\bf V}_{i})w({\bf V}_{i})\mid S_{i}=1\}=\mathbb{E}\{f({\bf V}_{i})\mid S_{i}=0\},\text{ any measurable }f.$
Han et al., (2021) considered this connection to construct an exponential tilting estimator of $w({\bf V}_{i})$. Relatedly, Josey et al., (2022); Chen et al., (2023) used entropy balancing of Hainmueller, (2012) to estimate $w({\bf V}_{i})$.
To account for the possible imbalance between the source and target samples (i.e., $n_{s}$ and $n_{t}$ may differ a lot) and to enable covariate balancing, we proceed with the entropy balancing method in (12) that falls into the second category. The solutions $\widehat{w}_{i}$ of entropy balancing are characterized in Lemma C.1.
Lemma C.1
The solution of (12) is $\widehat{w}_{i}=\exp(\widehat{\alpha}+\widehat{\boldsymbol{\beta}}^{\rm T}{\bf V}_{i})$, where $(\widehat{\alpha},\widehat{\boldsymbol{\beta}})$ is solution to
$\min_{\alpha,{\boldsymbol{\beta}}}\dfrac{1}{n_{s}}\sum_{i\in{\cal I}_{s}}\exp(\alpha+{\boldsymbol{\beta}}^{\rm T}{\bf V}_{i})-\alpha-\dfrac{1}{n_{t}}\sum_{i\in{\cal I}_{t}}{\boldsymbol{\beta}}^{\rm T}{\bf V}_{i}.$ (C.3)
Lemma C.1 is a special case of Proposition 1 of Chen et al., (2023). The dual problem (C.3) is an unconstrained convex optimization problem and numeric solutions can be efficiently solved by algorithms like the Newton-Raphson method. The implementation is performed using the optim function in R.
C.3 Proof of Theorem 4.2 and Theorem A.2
@@
We prove Theorem A.2, the EIF for a general outcome under sensitivity model (A.1). It includes Theorem 4.2 as a special case for a binary outcome. To simplify notation, we suppress the dependence of the TATE on $\Gamma_{a}$ and denote the expected potential outcome on the target population at treatment level $a$ as $\theta_{a}$ for $a=0,1$. We also drop the subscript $i$ and denote by ${\bf O}$ a generic random variable, which consists of $({\bf X},Y,S=1)$ for the source and $({\bf V},S=0)$ for the target. We recall that we have defined $\gamma_{a}={\textnormal{log}}(\Gamma_{a})$.
@@
We start with the case where $\pi({\bf X})$ is unknown and therefore considered as a nuisance parameter. For clarify we denote its true value as $\pi_{0}({\bf X})$. Denote by $p_{{\bf V}\mid S=1}$, $p_{{\bf X}\mid{\bf V},S=1}$, $p_{Y\mid{\bf X},A,S=1}$ the density functions of the conditional distributions of ${\bf V}\mid S$, ${\bf X}\mid{\bf V},S=1$ and $Y\mid{\bf X},A,S=1$, respectively. For a generic observation ${\bf O}$, the log-likelihood can be written as
@@
$\mathbb{E}\{\exp(\gamma_{1}Y)Y\mid{\bf X},A=1,S=1\}=$ $\Gamma_{1}\mu_{1}({\bf X}),$
$\mathbb{E}\{\exp(\gamma_{1}Y)\mid{\bf X},A=1,S=1\}=$ $\Gamma_{1}\mu_{1}({\bf X})+1-\mu_{1}({\bf X}),$
$\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})Y^{(1)}\mid{\bf V},S=1\}=$ $\Gamma_{1}\rho_{1}({\bf V}),$
$\mathbb{E}\{\exp(\gamma_{1}Y^{(1)})\mid{\bf V},S=1\}=$ $\Gamma_{1}\rho_{1}({\bf V})+1-\rho_{1}({\bf V}).$
@@
$=$ $\dfrac{S_{i}w({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{a}}{[\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\pi({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\pi({\bf X}_{i})}\right\}\{Y_{i}-\mu_{a}({\bf X}_{i})\}+\mu_{a}({\bf X}_{i})-\rho_{a}({\bf V}_{i})\right]$
$+\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{a}\rho_{a}({\bf V}_{i})}{\Gamma_{a}\rho_{a}({\bf V}_{i})+1-\rho_{a}({\bf V}_{i})},$
@@
$=$ $\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{a}}{[\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})]^{2}}\left[\left\{\dfrac{A_{i}}{\widehat{\pi}({\bf X}_{i})}+\dfrac{1-A_{i}}{1-\widehat{\pi}({\bf X}_{i})}\right\}\{Y_{i}-\widehat{\mu}_{a}({\bf X}_{i})\}+\widehat{\mu}_{a}({\bf X}_{i})-\widehat{\rho}_{a}({\bf V}_{i})\right]$
$+\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\dfrac{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})}{\Gamma_{a}\widehat{\rho}_{a}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})},$
@@
In particular, if $\Gamma_{1}=\gamma_{0}=1$, there exists a constant $C$ such that
@@
$=$ $\mathbb{E}\{\widehat{\varphi}_{1}({\bf O}_{i})\}-\mathbb{E}\left[\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right]$
$=$ $\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\dfrac{A}{\widehat{\pi}({\bf X}_{i})}\{\mu_{1}({\bf X}_{i})-\widehat{\mu}_{1}({\bf X}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{a}({\bf V}_{i})]^{2}}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{1-S_{i}}{\widehat{\mathbb{P}}(S_{i}=0)}\dfrac{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})}{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})}-\dfrac{1-S_{i}}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{1}\rho_{1}({\bf V}_{i})}{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})}\right]$
$=$ $\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\widehat{\mathbb{P}}(S_{i}=1)\widehat{\pi}({\bf X}_{i})}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\pi}({\bf X}_{i})-\pi({\bf X}_{i})\}\{\widehat{\mu}_{1}({\bf X}_{i})-\mu_{1}({\bf X}_{i})\}\right]$
$-\mathbb{E}\left[\left\{\dfrac{\mathbb{P}(S_{i}=1)-\widehat{\mathbb{P}}(S_{i}=1)}{\widehat{\mathbb{P}}(S_{i}=1)\mathbb{P}(S_{i}=1)}+\dfrac{1}{\mathbb{P}(S_{i}=1)}\right\}S_{i}\widehat{w}({\bf V}_{i})\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left(\dfrac{(1-S_{i})\Gamma_{1}\left[\widehat{\rho}_{1}({\bf V}_{i})\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}-\rho_{1}({\bf V}_{i})\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\right]}{\mathbb{P}(S_{i}=0)\widehat{\mathbb{P}}(S_{i}=0)\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right)$
@@
$-\mathbb{E}\left[\dfrac{S_{i}\widehat{w}({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{1}}{[\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})]^{2}}\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\}\right]$
$+\mathbb{E}\left[\dfrac{(1-S_{i})}{\mathbb{P}(S_{i}=0)}\dfrac{\Gamma_{1}}{\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}\right]$
@@
$+\mathbb{E}\left[\dfrac{S_{i}}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{1}\{1-\Gamma_{1}\}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}\left\{\widehat{w}({\bf V}_{i})-w({\bf V}_{i})\right\}}{\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}^{2}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right]$
$+\mathbb{E}\left[\dfrac{Sw({\bf V}_{i})}{\mathbb{P}(S_{i}=1)}\dfrac{\Gamma_{1}\{1-\Gamma_{1}\}\left\{\widehat{\rho}_{1}({\bf V}_{i})-\rho_{1}({\bf V}_{i})\right\}^{2}}{\{\Gamma_{1}\widehat{\rho}_{1}({\bf V}_{i})+1-\widehat{\rho}_{1}({\bf V}_{i})\}^{2}\{\Gamma_{1}\rho_{1}({\bf V}_{i})+1-\rho_{1}({\bf V}_{i})\}}\right]$
@@
When $\Gamma_{1}=0$, following the same procedure and using the fact that $\Gamma_{1}=1$, we have